OpenStudy (katherinesmith):

can someone please help me find and graph the focus and directrix of this equation.

OpenStudy (katherinesmith):

the equation is y = (1/32)x^2. i graphed it with my calculator and drew it on my paper. i don't understand how to find the focus and directrix.

OpenStudy (anonymous):

you need the vertex first

OpenStudy (katherinesmith):

how do i find the vertex.

OpenStudy (anonymous):

general form for this kind of problem is \[4p(y-k)=(x-h)^2\] but in your case you have \[y=\frac{1}{32}x^2\] lets make it look like the top one

OpenStudy (anonymous):

multiply both sides by \(32\) and get \[32y=x^2\] which we can visualize as \(32(y-0)=(x-0)^2\) making the vertex the origin \((0,0)\)

OpenStudy (anonymous):

satellite plz come help me when you done

OpenStudy (anonymous):

then we have \(4p=32\) so what is \(p\)?

OpenStudy (katherinesmith):

how did you even get 4p = 32 from that? like i literally don't understand.

OpenStudy (katherinesmith):

i know the focus is (0, p) and i guess you're saying p is 8 but i don't know how you managed to get 4p = 32.

OpenStudy (anonymous):

i multiplied both sides of \[y=\frac{1}{32}x^2\] by \(32\) to make it look more like \[4py=x^2\]

OpenStudy (katherinesmith):

well am i right in saying the focus is 0,8?

OpenStudy (anonymous):

this \(4p(y-k)=(x-h)^2\) is the general form of a parabola that opens up or down the vertex is \((h,k)\) and \(p\) is the distance from the vertex to the focus, and also to the directrix

OpenStudy (anonymous):

was that the answer satellite

OpenStudy (anonymous):

once we make it look like that , we read off \(4p=32\) and then we know \(p=8\)

OpenStudy (katherinesmith):

so the focus is 0,8 and the directrix is y = -p so the directrix is -8 on the y axis

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well not "-8 on the y axis" but rather the line \(y=-8\) which is what i assume you meant

OpenStudy (katherinesmith):

yeah. i understand how to find the focus and directrix but i still don't understand how to find p. like my next one is x = -(1/12)y^2, i know that this is a vertical directrix now instead of a horizontal but i still don't know how to find p. @satellite73